OpenStudy (anonymous):

What is the surface area of a conical grain storage tank that has a height of 43 meters and a diameter of 14 meters? Round the answer to the nearest square meter. (1 point) 1,100 square meters 1,112 square meters 2,507 square meters 2,605 square meters

OpenStudy (wolf1728):

Lateral Area of a Cone = (π • r • slant height) slant height² = height² + r² slant height² = 43² + 7² slant height = sq root (1,898) slant height = 43.566 Cone Area = 3.14159265 * 7 * 43.566 = 958.066 Bottom Area = PI * r² = 3.14159265 * 49 = 153.938 Total area = 1,112.004 square meters